OpenStudy (anonymous):

How many positive, three-digit integers contain at least one 3 as a digit but do not contain a 5 as a digit?

OpenStudy (anonymous):

What are the different classes? The total sample space, 100-999, is 900

OpenStudy (turingtest):

I almost never understand what you write moneybird, is that the answer?

OpenStudy (turingtest):

that makes sense...

OpenStudy (anonymous):

Oh wait I get it now. The number of 3 digit integers that do not contain either 3 or 5 as digits is 7*8*8=448. The number of 3 digit integers that just do not contain a 5 as a digit is 8*9*9=648. 648-448=200 :)

OpenStudy (turingtest):

yeah I think she is right, those 9's are wrong because they cannot be either 5 nor 3. If they could be three we would be counting some numbers twice, right?